Two balls are drawn in succession without replacement from an urn containing 5 red balls and 6 blue balls. Let Z be the random variable representing the number of Blue balls. Find the values of the random variable Z

1
Expert's answer
2022-03-04T05:10:25-0500

Let's denote R - red ball, B - blue ball.

Sample space S is all possible outcomes.


"S=\{RR, RB, BR, BB\}"

The possible values of the random variable "Z" are "0, 1, 2."

"\def\arraystretch{1.5}
 \begin{array}{c:c}
 Possible \ Outcomes & Z \\ \hline
 RR & 0 \\
 \hdashline
 RB & 1 \\
 \hdashline
 BR & 1 \\
 \hdashline
 BB & 2 \\
 \hdashline
\end{array}"


Construct the probability distribution of the random variable


"P(RR)=\dfrac{5}{11}(\dfrac{4}{10})=\dfrac{2}{11}"

"P(RB)=\dfrac{5}{11}(\dfrac{6}{10})=\dfrac{3}{11}"

"P(BR)=\dfrac{6}{11}(\dfrac{5}{10})=\dfrac{3}{11}"

"P(BB)=\dfrac{6}{11}(\dfrac{5}{10})=\dfrac{3}{11}"

"\def\arraystretch{1.5}
 \begin{array}{c:c}
 z & 0 & 1 & 2 \\ \hline
 p(z) & 2/11 & 6/11 & 3/11 
\end{array}"